OpenStudy (anonymous):

How to solve for x and y using elimination method? -4y-11x=36; 20=-10x-10y

OpenStudy (anonymous):

Im so confused on how to line the equations up

OpenStudy (ranga):

Rearrange the terms so that it has the x term first, then the y term and the constant term to the right. -4y-11x=36 can be rewritten as: -11x - 4y = 36 Multiply each term by -1 11x + 4y = -36 -------- (1) The next equation is: 20=-10x-10y which can be rewritten as: -10x - 10y = 20 This equation could be further simplified by dividing each term by -10: x + y = -2 -------- (2) Now can you solve for x and y?

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

Yeah now can! Thank you so much!

OpenStudy (ranga):

You are welcome.